# Application of Analytical Hierarchical Process and its Variants on Remote Sensing Datasets

### v1

@@ 3.5.2. Difference between 1-N AHP and AHP
3.5.3. Difference between each Fuzzy AHP case (Mean, Max 1, and Max 2) and AHP
@@
In the Mean Fuzzy AHP Case, we sampled some of the simulations, and most of them had a rank reversal between rainfall and drainage density, causing those layers to reduce the robustness of the results. In Max Case 1 and Max Case 2, both approaches were in agreement with each other across the latter half of the Ganga buffer, but we see that in the high-altitude areas they were tending in opposite directions. In Max Case 1 the AHP suggested a higher vulnerability than Fuzzy AHP, primarily affected by Population Density, and in Max Case 2, Fuzzy AHP returned a higher value for vulnerability scores, owing to the Annual Rainfall variable.
3.5.4. Difference between Population density and AHP
@@
The differences between AHP and the Population density layer post-standardization were the highest among all approaches, and there were many areas where there is a non-zero value for this layer. This indicates that AHP is amply influenced by the other parameters and proves the effectiveness of the AHP approaches for creating a good composite variable.
3.5.5. Analysis of Variances
@@
To compare the predictions for multiple variants at once, we calculated the standard deviation between three approaches: AHP, ANP, Nested AHP for each pixel and visualized them for our region of interest (Figure 13). We noticed that the deviations were high for the tier-two cities Prayagraj/Allahabad and Varanasi. Varanasi had the highest vulnerability in terms of AHP, but the other variants of AHP brought more emphasis to Prayagraj/Allahabad as well. There were also some high deviations in the high-altitude regions near Rishikesh - a high-population city in Uttarakhand, suggesting disagreements between the approaches.
Figure 14. Normalized AHP and its Standard Deviation
@@
Figure 13. Standard deviations of Normalized layers for the approaches AHP, ANP, Nested AHP, where darker values suggest lower standard deviation, and the lighter values suggest high-standard deviations.
3.6. Composite Layer for all Approaches
@@
To create a composite layer to capture the effect of each variant of AHP across the river buffer, we took a normalized base layer for AHP, and overlaid the outliers (¿=2 standard deviations) for each of the difference layers for Fuzzy AHP, Nested AHP, and ANP (Figure 14). To understand the distribution of the AHP base layer, and the difference layers for Fuzzy AHP, Nested AHP, and ANP, we also created the histograms of each layer (Figure 15). In the composite layer, the primary colors (red, blue, yellow) were used to depict the outliers for the individual layers, and secondary colors (orange, green, purple) were used when the effects of two AHP variants were combined. Given that the outliers could be both positive and negative, we also created separate maps for the two. These composite layers can be used as a way to depict a single variable for river Ganga’s vulnerability with the effect of all AHP variants. Through these composite layers, we are illustrating the effect of AHP relative to the unique effects of the variants wherever they disagree with AHP.
@@
We also selected four important subsections of the river buffer where the composite variable was effective in capturing the effects of each AHP variation (AHP, Nested AHP, and ANP). There are some small areas near the river body which had all three outliers. We can see the effect of the combination of two variants in Figure 16 subsection C, where there were a lot of outliers for both Nested AHP+ANP (violet) and Nested AHP+Fuzzy AHP (green) variants.
Figure 15. Composite Layers Outliers
@@
Figure 14. The composite layer a.) that can be used to describe the vulnerability of the river Ganga. We assign different colors for the outliers of each variant, and to showcase the positive and negative outliers, we split them into two charts b.) and c.)
Figure 16. Normalized AHP & its Difference with Normalized AHP Variants
@@
Figure 15. Figure 15a. shows the Normalized AHP layer with its standard deviation, while the rest of the figures show the difference between Normalized AHP and its corresponding AHP Variants: ANP, Fuzzy AHP, and Nested AHP. Figure 15b. is Normalized AHP minus Normalized ANP. Figure 15c. is Normalized AHP minus Normalized Fuzzy AHP. Figure 15d. is Normalized AHP minus Normalized Nested AHP
Figure 17. Focused Composite Layers Outliers
@@
Figure 16. Subsection analysis of the composite variable for four subsections A, B, C, D. To cover a range of distributions, we considered two rural (A, D), and two urban (B, C) subsections. Out of these, Subsection B has the city Kanpur in them, while subsection C is near Prayagraj/Allahabad and Varanasi, highly populated areas along the Ganga.
4. Discussion
@@
Additional data layers such as GIS-based applications that track impacts may be integrated into the data process to support and hone vulnerability assessments and thus AHP factor weights as well as improve output resolution and validation. For instance, acute factors may be known unknowns (e.g., mining locations) and unknown/ill defined unknowns (e.g., location and breadth of endangered species habitat). Similarly, potential widespread impacts such as loss of biodiversity and fragmentation of biodiversity richness are chronic factors that were excluded from our AHP vulnerability decision space but may be included in future efforts. Indeed, the approach of sand mining and waste disposal tracked by Bayazidy et al. (2024)(2, ) may be used to identify specific instances of these and other acute factors that require precise and timely identification. In addition, coupling biodiversity and the conservation imperative analysis of Dinerstein et al. (2024)(5, ) with our vulnerability assessment can further prioritize areas for conservation that were scored as low biodiversity and extremely low vulnerability. Moreover, emerging factors may be added as data are collected and the breadth and depth of the decision space grow. For instance, the environmental spread of antibiotic resistance in freshwater is tracked with molecular genetic techniques (Yang et al., 2018(29, ); Liu et al., 2023(11, )), and coupling these data with their spatial locations can add antibiotic resistance as a vulnerability factor. Finally, high-resolution data from UAV such as that produced by Tripathi et al. (2024)(27, ) during their effort to combine remote sensing with drone data while cataloging riparian vegetation along the Ganga can augment and further refine LULC information to better refine vulnerability scores.
@@
Ultimately, networks of models may be brought together to enrich landscape digital twins for design, decision making, and monitoring of real world features of landscapes. These twins become digital representations of natural and built features in landscapes. For river systems, digital twins may consist of riverscapes, the concept that integrates longitudinal characteristics of rivers with their land features (Zhou et al., 2014)(31, ). Indeed, the basis for such twins already exists in regions prone to flooding. For instance, Nested AHP was used to predict five categories of flood susceptibility with nine factors based on remote sensing data in addition to soil texture, geomorphology, and geology (Das and Gupta, 2021)(4, ). Furthermore, AHP with nesting of four factors (land use, percent green space, per capita sewer length, and slope) was used to produce a flood vulnerability index in a section of Hanoi, Vietnam (Loi, 2023)(10, ), and a flood susceptibility map for all of Hanoi was made with a 9-factor AHP that was ranked by subject matter experts (Nguyen et al., 2024)(14, ). AHP models that include cultural, socioeconomic and environmental criteria provide for wider assessments of the human condition. Recently, 20 factors representing these criteria as well as security and service functions were placed in a hierarchical structure and assessed through a survey of subject matter experts to realize a quality of life score (Withanage et al., 2024)(28, ). Of note, they detected the inverse of our vulnerability scores in that quality of life tended to decrease away from city centers and the services they provide whereas cities consistently were sources of high vulnerability to the river Ganga. Thus, the perspective of the modeling scenario must be understood and integrated into the larger decision space of balance between human endeavors and nature. A richness of alternative perspectives can drive sustainable development so that factors that improve quality of life can be implemented with policies that recognize and mitigate their accompanying sources of vulnerabilities.
5. Conclusion
@@
In this study, we assessed the vulnerability of the river Ganga to pollution along a 20 km wide corridor covering 26,609.4 km2 of riverscape. We used AHP to perform a dimensionality reduction of six remote sensing datasets to create a single variable for river vulnerability. To mitigate structural limitations to AHP, our approach brought together a suite of methods and produced a set of comparative metrics to identify and refine vulnerability scores. Of note, urban landscape features with extremely high vulnerability scores and those at the interface with vulnerable areas were identified. These findings provide the basis to rational prioritization of pollution mitigation and a list of locations for future rejuvenation. Moreover, sites currently with low vulnerability scores may be targeted for conservation or sustainable development practices to prevent their degradation. Our vulnerability score, together with other metrics such as a quality of life score (Withanage et al., 2024)(28, ), can enrich digital twins beyond simplistic animations and reveal their underlying environmental, social, and economic characteristics.
6. CRediT Authorship
Sarthak Arora: Conceptualization, Methodology, Software, Formal analysis, Data curation, Visualization, Investigation, Validation, Writing- Reviewing and Editing, Writing- Original draft preparation
Michael Warner: Data curation, Software, Visualization, Writing- Original draft preparation, Writing- Reviewing and Editing
Ariel Chamberlain: Visualization, Formal analysis, Data Curation, Writing- Reviewing and Editing
James C. Smoot: Supervision, Conceptualization, Project Administration, Writing- Reviewing and Editing, Writing- Original draft preparation
Nikhil Raj Deep: Conceptualization, Software, Writing- Reviewing and Editing
Claire Gorman: Conceptualization, Writing- Reviewing and Editing
Anthony Acciavatti: Writing- Reviewing and Editing, Supervision, Conceptualization, Project Administration
7. Declaration of competing interest
@@
8. Acknowledgements
The authors thank Markley Boyle and Sarah Bergmann for thoughtful discussions throughout the project. We also thank Aaron Hirsh for organizing Collaborative Earth.
@@ 8. Acknowledgements
Table S1. Shortcomings of AHP analysis rubric.
@@ 5.2.10 Quantitative Data N
5.2.13 Relative Importance Between Criteria Is Considered Constant Y This is the crux of Steele et al. (2009) (S_Steele, ) concern with respect to relative normalization and absolute normalization which boils down to the decision space is unlikely to be completely known. 1-N AHP
@@ 5.2.15 A Preference and the Peculiar Meaning of Its Inverse Value N
Table S2. Shortcomings of AHP analysis rubric (continued).
@@ 5.2.29 Problem structuring N
Figure S1. AHP Box and Whisker Plot Summary
Figure S1. Panels (a) through (e) are box and whisker plots for the length of the Ganga contained in the area of analysis. Data distribution was tested and results are given for each factor label. Whiskers show the range of values, the box “shoulders” are the inter-quartile range, and the median is shown as a band.
Figure S2. AHP Histogram Summary
Figure S3. AHP Histogram Summary (continued)
Figure S2. Panels (a) through (l) are the histograms and log-histograms to assess the distributions of the AHP factors along the length of the Ganga contained in the area of analysis.
Table S3. Data Sources
@@
NASADEM NASA NASADEM Digital Elevation 30m (3 bands) Link NASA JPL, 2020 (S_NASA, )
WorldCover ESA WorldCover 10m v200 Link Zanaga, 2021 (S_Zanaga, )
GPWv411 UN-Adjusted Population Density (Gridded Population of the World Version 4.11) Link CIESIN, 2018 (S_CIESIN, )
CHIRPS Daily Climate Hazards Group InfraRed Precipitation With Station Data (Version 2.0 Final) Link Funk, 2015 (S_Funk, )
Landsat USGS Landsat 8 Level 2, Collection 2, Tier 1 Link USGS (S_USGS, )
@@
The importance of factors influencing naala vulnerability were placed in the following order: Population Density (PD) ¿ LULC ¿ Rainfall (RAIN) ¿ Drainage Density (DD) ¿ Slope ¿ LST, and the factor importance rankings were inputs into an AHP pairwise comparison matrix (Supplemental Table. S3). In some situations, AHP importance rankings may produce inconsistent outcomes due to matrix properties. The Consistency Index (S1) and Consistency Ratio (Consistency Index/Random Index) were developed to test for inconsistency with a Consistency Ratio ¡ 0.1 indicating the rankings were consistent (Saaty, 1980)(S_Saaty, ).
(S1) $CI=\frac{\lambda_{max}-n}{n-1}$
Supplemental Table. S3a is the pairwise comparison matrix; Supplemental Table. S3b is the Normalized AHP pairwise comparison matrix. The Principal Eigenvalue is the mean of the eigenvectors, and n is the number of factors - six.
Table S4. AHP Pairwise and Normalized Comparison Matrixa
a.) AHP pairwise comparison matrix PD LULC RAIN DD SLOPE LST PD 1 2 4 5 7 9 LULC 0.5 1 3 4 5 7 RAIN 0.25 0.33 1 2 4 5 DD 0.2 0.25 0.5 1 3 5 SLOPE 0.14 0.2 0.25 0.33 1 4 LST 0.11 0.14 0.2 0.2 0.25 1 Total 2.2 3.92 8.95 12.53 20.25 31
@@
Data for each AHP factor were divided into five vulnerability classes using the Jenks natural breaks method (Jenks and Caspall, 1971)(S_Jenks, ). The Jenks method is widely used within GIS packages to generate variance-minimization classification. Breaks were typically uneven, and were selected to separate values where large changes in value occur. For this, we used the boxplots and histograms for the layers to identify the natural breaks for each factor (Supplemental Figure S1 and Supplemental Figure S2).
@@
Minimum and maximum values for the five factor classes of population density, slope, rainfall, and drainage density were averaged, normalized, transformed to a 9-point scale to create the factor rankings, and fit to the Saaty AHP scale based on their relationships with pollution production or its transport as well as their breadth across a 9-point scale (Table 2 and Supplemental Table S4). The relationship between population density and pollution production was assumed linear based on municipal solid waste production in India (Goel, 2008)(S_Goel, ), and the factors were ranked as 1, 4, 6, 8, and 9. Rainfall factor rankings spanned 1 through 5 on the Saaty scale because the range between lowest and highest average annual precipitation was a factor of 2 (1, 3, 4, 5, 5). Drainage density can have a nonlinear relationship with pollution transport (Bedient et al. 1978)(S_Bedient, ), and although the range of drainage density values was more than 20-fold, its factors were ranked as 1, 6, 8, 9, and 9. The relationship between slope and transport is nonlinear for slopes greater than  26 degrees, and linear for slopes less than 26 degrees (Andrews and Bucknam, 1987)(S_Andrews, ). So slope factor rankings were ranked 1, 1.5, 3, 6, and 9. Minimum and maximum temperature data per factor rank were used to calculate normalized temperature dependent microbial growth rates with a Q10 (temperature coefficient) of 1.75. Then the growth rates were transformed to a 9-point scale and scored on the Saaty scale (1, 4, 5, 6, 8). LULC factor ranks were based on ESA WoldCover map (Zanaga, 2021(S_Zanaga, )) categories (built-up, cropland, bare/sparsely vegetated, tree cover/grassland, water/wetlands) (1,3,4,5,7). Built-up was treated as the factor rank 1 because of the breadth of pollution sources (e.g., municipal solid wastes, pathogens, excess nutrients, and heavy metals) and amount of impervious surfaces that lead to pollution transport. Cropland was ranked 3 due to runoff of fertilizers and agrochemicals. Bare/sparsely vegetated land was ranked 4 due to moderate transport risk. Tree cover and grassland were ranked 5 on the Saaty scale because these categories represent natural ecosystems, and water/wetlands were ranked as 7 on the Saaty scale, because Ganga was the point of pollution deposition.
Table S5. Nested AHP: Factor Vulnerability Rankings for all Factors
@@
Barren/ 0.25 1 1 1 3
Sparse
Vegetation
Trees/ Grass/ 0.2 0.5 1 1 2
Shrubs
Water / 0.14 0.25 0.33 0.5 1
Wetland
@@
Table S6. Continued: Nested AHP: Factor Vulnerability Rankings for all Factors
Slope (Angle %) ¿35 24-35 45650 45394 ¡4
¿35 1 1.5 3 6 9
@@
Pairwise comparisons were made to describe the relative importance of the ANP criteria (AHP factors) on influencing each other (Supplemental Figure S3). Criteria that did not contribute to an inner dependence matrix were scored with a zero. When identifying the relative importance of Land Use Land Cover (LULC) compared to other criteria on influencing Population Density (PD), it was scored with a 2 because of the importance of built-up and agriculture categories in LULC and their direct relationships with population densities. Rainfall and slope were scored with a 5 because historically there were natural limits to both that influenced urbanization and agriculture although this is changing (Zhou et al., 2021(S_Zhou, ); Chen et al., 2022(S_Chen, ); Shi et al., 2022(S_Shi, ); Wang et al., 2023(S_Wang, ); Ullah et al., 2023(S_Ullah, )). Drainage Density (DD) was scored an 8 because the natural drainage density is modified by people to meet agricultural and societal requirements suggesting it has a minor influence on PD. When identifying the relative importance of PD compared to other criteria on influencing LULC, it was scored with a 2 as a reciprocal of the LULC-PD inner dependence. Rainfall and temperature are major drivers of land cover type, and as such each was scored with a 4. Slope and DD shape landscapes within land cover types, and each was scored with a 7. When identifying the relative importance of LULC compared to other criteria on influencing rainfall, it was scored with a 7 and PD was scored with an 8 because of the potential for built-up land to influence local precipitation (Liu and Niyogi, 2019)(S_Liu, ). Slope may influence the speed of runoff during rainfall and was scored with a 5. When identifying the relative importance of criteria on influencing DD, PD was scored a 2, LULC was scored a 4, and slope was scored a 7, because DD is the product of both anthropogenic and natural influences. When identifying the relative importance of criteria on influencing temperature, PD was scored a 5, and LULC was scored an 8 to account for potential urban heat island effects.
Figure S4. ANP Parameter Inner Dependencies
Figure S3. Inner dependencies between ANP criteria (AHP factors) are shown as ovals. Arrows about the ovals indicate influence by a criterion on other criteria to its right and arrows below the ovals indicate influence by a criterion on other criteria to its left.
Figure S5. ANP with original AHP submatrix
Figure S4. These figures show the ANP’s vulnerability score with the original AHP submatrix; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Figure S4a. shows the exact locations and values, while figure S4b. shows the count for each bin range of the vulnerability in histogram form.

### v2

@@ 3.5.2 Difference between 1-N AHP and AHP
3.5.3 Difference between Fuzzy AHP cases (Mean, Case 1, and Case 2) and AHP
@@
In the Mean Fuzzy AHP Case, we sampled some of the simulations, and most of them had a rank reversal between rainfall and drainage density, causing those layers to reduce the robustness of the results. In Case 1 and Case 2, both approaches were in agreement with each other across the latter half of the Ganga buffer, but we see that in the high-altitude areas they were tending in opposite directions. In Case 1 the AHP suggested a higher vulnerability than Fuzzy AHP, primarily affected by Population Density, and in Case 2, Fuzzy AHP returned a higher value for vulnerability scores, owing to the Annual Rainfall variable.
3.5.4 Difference between Population density and AHP
@@
The differences between the standardized values of AHP and the Population density layers were the highest among all approaches, and there were many areas where there is a non-zero value for this layer. This indicates that AHP is amply influenced by the other parameters and proves the effectiveness of the AHP approaches for creating a good composite variable.
3.5.5 Analysis of Variances
@@
To compare the predictions for multiple variants at once, we calculated the standard deviation between three approaches: AHP, ANP, Nested AHP for each pixel and visualized them for our region of interest (Fig 13). We noticed that the deviations were high for the tier-two cities Prayagraj/Allahabad and Varanasi. Varanasi had the highest vulnerability in terms of AHP, but the other variants of AHP brought more emphasis to Prayagraj/Allahabad as well. There were also some high deviations in the high-altitude regions near Rishikesh - a high-population city in Uttarakhand, suggesting disagreements between the approaches.
@@
Fig 13. Standard deviations of Normalized layers for the approaches AHP, ANP, Nested AHP, where darker values suggest lower standard deviation, and the lighter values suggest high-standard deviations.
3.6 Composite Layer for all Approaches
@@
To create a composite layer to capture the effect of each variant of AHP across the river buffer, we took a normalized base layer for AHP, and overlaid the outliers (>=2 standard deviations) for each of the difference layers for Mean Fuzzy AHP, Nested AHP, and ANP (Fig 14). To understand the distribution of the AHP base layer, and the difference layers for Mean Fuzzy AHP, Nested AHP, and ANP, we also created the histograms of each layer (Fig 15). In the composite layer, the primary colors (red, blue, yellow) were used to depict the outliers for the individual layers, and secondary colors (orange, green, purple) were used when the effects of two AHP variants were combined. Given that the outliers could be both positive and negative, we also created separate maps for the two. These composite layers can be used as a way to depict a single variable for river Ganga’s vulnerability with the effect of all AHP variants. Through these composite layers, we are illustrating the effect of AHP relative to the unique effects of the variants wherever they disagree with AHP.
@@
We also selected four important subsections of the river buffer where the composite variable was effective in capturing the effects of each AHP variation (AHP, Nested AHP, and ANP). There are some small areas near the river body which had all three outliers. We can see the effect of the combination of two variants in Fig 16 subsection C, where there were a lot of outliers for both Nested AHP + ANP (violet) and Nested AHP + Mean Fuzzy AHP (green) variants.
@@
Fig 14. The composite layer (a) that can be used to describe the vulnerability of the river Ganga. We assign different colors for the outliers of each variant, and to showcase the positive and negative outliers, we split them into two charts (b) and (c)
@@
Fig 15. Here, (a) shows the Normalized AHP layer with its standard deviation, while the rest of the figures show the difference between Normalized AHP and its corresponding AHP Variants: ANP, Mean Fuzzy AHP, and Nested AHP. (b) is Normalized AHP minus Normalized ANP. (c) is Normalized AHP minus Normalized Mean Fuzzy AHP. (d) is Normalized AHP minus Normalized Nested AHP
@@
Fig 16. Subsection analysis of the composite variable for four subsections A, B, C, D, with different panels numbered 1 (all outliers), 2 (positive outliers), and 3 (negative outliers). To cover a range of distributions, we considered two rural (A, D), and two urban (B, C) subsections. Out of these, Subsection B has the city Kanpur in them, while subsection C is near Prayagraj/Allahabad and Varanasi, highly populated areas along the Ganga.
4 Discussion
@@
Additional data layers such as GIS-based applications that track impacts may be integrated into the data process to support and hone vulnerability assessments and thus AHP factor weights as well as improve output resolution and validation. For instance, acute factors may be known unknowns (e.g., mining locations) and unknown/ill defined unknowns (e.g., location and breadth of endangered species habitat). Similarly, potential widespread impacts such as loss of biodiversity and fragmentation of biodiversity richness are chronic factors that were excluded from our AHP vulnerability decision space but may be included in future efforts. Indeed, the approach of sand mining and waste disposal tracked by Bayazidy et al. (2024) may be used to identify specific instances of these and other acute factors that require precise and timely identification. In addition, coupling biodiversity and the conservation imperative analysis of Dinerstein et al. (2024) with our vulnerability assessment can further prioritize areas for conservation that were scored as low biodiversity and extremely low vulnerability. Moreover, emerging factors may be added as data are collected and the breadth and depth of the decision space grow. For instance, the environmental spread of antibiotic resistance in freshwater is tracked with molecular genetic techniques (Yang et al., 2018; Loi, 2023), and coupling these data with their spatial locations can add antibiotic resistance as a vulnerability factor. Finally, high-resolution data from UAV such as that produced by Tripathi et al. (2024) during their effort to combine remote sensing with drone data while cataloging riparian vegetation along the Ganga can augment and further refine LULC information to better refine vulnerability scores.
@@
Ultimately, networks of models may be brought together to enrich landscape digital twins for design, decision making, and monitoring of real world features of landscapes. These twins become digital representations of natural and built features in landscapes. For river systems, digital twins may consist of riverscapes, the concept that integrates longitudinal characteristics of rivers with their land features (Zhou et al., 2014). Indeed, the basis for such twins already exists in regions prone to flooding. For instance, Nested AHP was used to predict five categories of flood susceptibility with nine factors based on remote sensing data in addition to soil texture, geomorphology, and geology (Das and Gupta, 2021). Furthermore, AHP with nesting of four factors (land use, percent green space, per capita sewer length, and slope) was used to produce a flood vulnerability index in a section of Hanoi, Vietnam (Liyanage and Yamada, 2017), and a flood susceptibility map for all of Hanoi was made with a 9-factor AHP that was ranked by subject matter experts (Nguyen et al., 2024). AHP models that include cultural, socioeconomic and environmental criteria provide for wider assessments of the human condition. Recently, 20 factors representing these criteria as well as security and service functions were placed in a hierarchical structure and assessed through a survey of subject matter experts to realize a quality of life score (Withanage et al., 2024). Of note, they detected the inverse of our vulnerability scores in that quality of life tended to decrease away from city centers and the services they provide whereas cities consistently were sources of high vulnerability to the river Ganga. Thus, the perspective of the modeling scenario must be understood and integrated into the larger decision space of balance between human endeavors and nature. A richness of alternative perspectives can drive sustainable development so that factors that improve quality of life can be implemented with policies that recognize and mitigate their accompanying sources of vulnerabilities.
5 Conclusion
@@
In this study, we assessed the vulnerability of the river Ganga to pollution along a 20 km wide corridor covering 26,609.4 km2 of riverscape. We used AHP to perform a dimensionality reduction of six remote sensing datasets to create a single variable for river vulnerability. To mitigate structural limitations to AHP, our approach brought together a suite of methods and produced a set of comparative metrics to identify and refine vulnerability scores. Of note, urban landscape features with extremely high vulnerability scores and those at the interface with vulnerable areas were identified. These findings provide the basis to rational prioritization of pollution mitigation and a list of locations for future rejuvenation. Moreover, sites currently with low vulnerability scores may be targeted for conservation or sustainable development practices to prevent their degradation. Our vulnerability score, together with other metrics such as a quality of life score (Withanage et al., 2024), can enrich digital twins beyond simplistic animations and reveal their underlying environmental, social, and economic characteristics.
6 Declaration of competing interest
@@
7 Acknowledgements
The authors thank Markley Boyer and Sarah Bergmann for thoughtful discussions throughout the project. We also thank Aaron Hirsh for organizing Collaborative Earth.
[…]
@@ 7 Acknowledgements
Table S1: Shortcomings of AHP analysis rubric.
@@ 5.2.10 Quantitative Data N
[…]
@@ 5.2.15 A Preference and the Peculiar Meaning of Its Inverse Value N
Table S2: Shortcomings of AHP analysis rubric (continued).
@@ 5.2.29 Problem structuring N
Fig S1. Panels (a) through (e) are box and whisker plots for the AHP factors Population Density (a), Annual Rainfall (b), Drainage Density (c), Slope (d), and Land Surface Temperature (e) for the length of the Ganga contained in the area of analysis. Data distribution was tested and results are given for each factor label. Whiskers show the range of values, the box "shoulders" are the inter-quartile range, and the median is shown as a band.
Fig S2. Panels (a) through (l) are the histograms and log-histograms to assess the distributions of the AHP factors along the length of the Ganga contained in the area of analysis. Panels (a), (c), (e), (g), (i) represent the histograms for Population Density, Annual Rainfall, Drainage Density, Slope, and Land Surface Temperature, respectively, while panels (b), (d), (f), (h), (j) respectively represent the log-histograms for the same. Panels (a) through (j) are binned using deciles, and coloured green to red for low to high vulnerability. (k) and (l) represent LULC and are color coded to their respective Land Use classes.
Table S3: Data Sources
@@
[…]
@@
[…]
Table S4: AHP Pairwise and Normalized Comparison Matrixa
a.) AHP pairwise comparison matrix
PD LULC RAIN DD SLOPE LST
PD 1 2 4 5 7 9
LULC 0.5 1 3 4 5 7
RAIN 0.25 0.33 1 2 4 5
DD 0.2 0.25 0.5 1 3 5
SLOPE 0.14 0.2 0.25 0.33 1 4
LST 0.11 0.14 0.2 0.2 0.25 1
Total 2.2 3.92 8.95 12.53 20.25 31
@@
[…]
@@
[…]
Table S5: Nested AHP: Factor Vulnerability Rankings for all Factors
@@
Barren/Sparse Vegetation 0.25 1 1 1 3
Trees/ Grass/ Shrubs 0.2 0.5 1 1 2
Water / Wetland 0.14 0.25 0.33 0.5 1
@@
Slope (Angle %) >35 24-35 45650 45394 <4
>35 1 1.5 3 6 9
@@
[…]
Fig S3. Inner dependencies between ANP criteria (AHP factors) are shown as ovals. Arrows about the ovals indicate influence by a criterion on other criteria to its right and arrows below the ovals indicate influence by a criterion on other criteria to its left.
Fig S4. These figures show the ANP’s vulnerability score with the original AHP submatrix; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
[…]